# From Pen to Prompt: How Creative Writers Integrate AI into their Writing Practice

### v2

@@ 6. Limitations
Our interview structure was shaped by the need to understand writers’ usual workflows, though we recognized that some participants might feel uncomfortable being observed while engaging in creative work. To mitigate this, participants could prerecord writing sessions or opt for a diary study. The first author also offered muted video during live sessions. Overall, most participants reported comfort and enjoyment in verbalizing and sharing their process. A minority noted heightened self-awareness or rushed pacing (e.g., one participant spent less time between iterations). Most felt that the sessions reflected their typical practice or explained any minor deviations.
@@ Appendix A Interview Protocol
[…]
@@
Follow-up Questions (25 min) In this section, our follow-up questions focused on the role of AI tools in the different stages of writing. Our aim was not only to understand how these tools are being used but also to explore the reasons behind the non-use of certain AI functionalities, which might reveal underlying value conflicts. Specifically, we delved into their experiences of writing a creative text with AI assistance. While the live writing session provided insights into the detailed interactions during the actual writing of text, these conversations allowed us to explore the role of AI in the planning stage of writing, such as the planning and development of character personas and story plots. We also asked participants to share their perspectives on the value of AI written text, AI attribution in writing, and any ethical concerns.
@@ Appendix B Participants’ AI use
These are the recorded answers from participants’ pre-interview survey answering the question ”Briefly, how do you use AI in your practice?”
@@ 17 Brainstorm ideas, mitigate writing block
Appendix C Writers’ Values
We identify four core values that creative writers want to preserve while using AI: authenticity, ownership & control, creativity, and craftsmanship. The descriptions provided for each value aim to distill what writers mean when they refer to these concepts—while noting that participants often interpreted these values in different ways.
@@ Appendix C Writers’ Values
Authenticity Importance of writing feeling that the writing comes from the author. (After AI generating a few lines) But beyond that, that’s when I start getting a little bit uncomfortable with how much am I losing myself in this writing process to something that’s not me authentically.” —P13 “Keeping it still sounding like me is, like, really important to me.” —P4
@@
Table 3. Writers’ values codes and example quotes
Appendix D Writers’ Relationships with AI
@@
Table 4. How writers see AI’s role in the writing process.
Appendix E Writers’ Specific Uses of AI

### v4

@@ 6. Limitations
Our interview structure was shaped by the need to understand writers’ usual workflows, though we recognized that some participants might feel uncomfortable being observed while engaging in creative work. To mitigate this, participants could prerecord writing sessions or opt for a diary study. The first author also offered muted video during writing sessions. Overall, most participants reported comfort and enjoyment in verbalizing and sharing their process. A minority noted heightened self-awareness or rushed pacing (e.g., one participant spent less time between iterations). Most felt that the sessions reflected their typical practice or explained how they might have done the writing slightly differently.
@@ Appendix A Interview Protocol
Here we outline the questions that guided our interviews and the live-writing observation sessions. The order of questions asked was also adjusted depending on what topics arose naturally through conversation, or skipped if not relevant. In line with grounded theory methodology, the questions evolved over the course of the study in response to participant behavior, emerging themes, and the dynamics of individual conversations. We conducted two rounds of updates to the interview guide to elicit richer, more relevant material as new insights and patterns emerged. Many of these questions came up naturally through the course of the interview and then were officially added to the interview guide. These updated questions are marked with (R1) and (R2).
General Writing Background: We asked a set of background and onboarding questions regarding participants’ writing practices and AI assistance in general.
  • What kind of writing do you currently do, and for how long?
  • What kind of formal or informal education have you had as a writer?
  • What motivates you to write? Which aspect of your writing experience do you enjoy the most?
  • How do you define success for your writing? (R1)
AI Background and Use:
  • What AI-powered writing platforms have you used?
  • What AI-powered writing do you currently use and how do you use them?
  • How did you begin using these AI platforms? How did you learn to use them? (R1)
  • What do you mainly use them for, and at what stage of your writing?
  • When do you reach out for AI? At which stages? What are your goals for writing with AI and why?
  • Which parts of using AI do you enjoy, and which parts do you dislike? (R1)
  • How would you describe your core values as a writer? What is important to you in the writing process? Which of those feel supported or hindered by AI collaboration?
Live-writing Session: Participants were invited to live-write a creative piece using AI in their unique workflows while sharing their screens. Writers were encouraged to talk through the process but ultimately do what felt natural to them. During the live-write, we would occasionally ask questions about motivations and feelings, giving space for the writers to reflect and evaluate their interactions with AI. We began by asking for context on what they were planning to write and what their goals were.
  • Why did you use AI here? / Why did you choose not to use it?
  • How did you come up with that prompt?
  • How are you evaluating the AI’s output?
  • What were you hoping to get from the AI here?
  • How do you feel about the results you got?
  • What is the reason for doing X (action participant performed on screen)?
  • Does it still feel like your idea/why do you feel comfortable using or not using the output?
  • Does this feel like the final text? Will you come back and edit it?
  • How strong of a plot or idea do you have when prompting the AI? Are you looking for ideas or looking to verbalize the vision you already have?
  • After the session:
    • How do you feel about the text that they wrote?
    • How do you see attribution and ownership?
    • Did you pick up on patterns or processes that you didn’t notice before? (R1)
Walkthrough of Previous Writing: Participants were also invited to share a piece (or pieces) of writing they had previously finished with the assistance of AI writing tools. We asked questions about what they remembered about the steps of the process as well as reflections of their views on creativity, attribution and ownership, and important moments from integrating AI into their workflows.
  • How long ago was this piece written? Do you remember how you used the AI?
    • Which ideas were from the AI’s suggestions?
    • Which words/phrases came from the AI’s suggestions?
  • Were there any memorable moments while writing this?
  • What were some moments that you remember the AI being helpful or frustrating?
  • How do you think about ownership and attribution in this piece?
@@
Workflow Reflections: In this section, our follow-up questions focused on the role of AI tools in the different stages of writing. Our aim was not only to understand how these tools are being used but also to explore the reasons behind the non-use of certain AI functionalities, which might reveal underlying value conflicts. Specifically, we dived into their experiences of writing a creative text with AI assistance. While the live writing session provided insights into the detailed interactions during the actual writing of text, these conversations allowed us to explore the role of AI in the planning stage of writing, such as the planning and development of character personas and story plots. We also asked participants to share their perspectives on the value of AI written text, AI attribution in writing, and any ethical concerns.
  • How would you describe your relationship with the AI or the role AI plays in your writing? Is there a point at which you would consider AI a co-author or collaborator? (R1)
  • How has your relationship with AI changed over time? (R2)
  • How do you think about disclosing your use of AI? (R1)
  • What is your understanding of how these AI tools work?
    • Has your view on AI changed over time as you’ve gotten more experience with it? (R2)
    • Have your views on originality, creativity, and ownership evolved through working with AI?
  • Does your AI use vary across genres? (R1)
  • What does an ideal future look like for AI and creative writing? (R1)
  • Do you have any concerns about using AI?
@@ Appendix B Participants’ AI use
These are the recorded answers from participants’ pre-interview survey answering the question ”Briefly, how do you use AI in your practice?” (Table 2).
@@ 17 Brainstorm ideas, mitigate writing block
Appendix C Summary of AI Platforms
Here we provide an overview of the most common AI writing platforms that participants used in this study. For each platform, we highlight the distinctive interface elements and main features (not exhaustive) that participants used in their creative process.
C.1. General Purpose Conversational AI
ChatGPT (Figure 4), Claude, Copilot, Perplexity and Gemini are all general purpose conversational AI platforms (not made for creative writing specifically). All use a chat-based interface where users type prompts or upload documents and receive text responses in a back and forth manner.
Figure 4. ChatGPT interface showing a chat interaction.
[…]
Screenshot of ChatGPT interface. The user asks ’What is a metaphor connecting bread and moss?’ ChatGPT responds below that with an initial metaphor ’Bread is moss domesticated.’ The interface shows a text input box at bottom.
C.2. Specialized Creative Writing Platforms
Unlike general-purpose conversational AI, several platforms are designed specifically for creative writers with features tailored to fiction writing, poetry, and other creative genres. These specialized platforms (NovelCrafter, NovelAI, Sudowrite, LAIKA, and Rexy) typically offer inline text generation that integrates directly with the writing process , along with custom AI models and unique features for different aspects of the writing process. These platforms combine AI assistance with organizational tools that writers can use independently of the AI features, such as support developing story plans and keeping track of plot points and characters. Below we summarize the common AI features present in many of these platforms.
  • Inline text generation: Writers can place their cursor in the middle of a paragraph, and the AI will generate a continuation directly in place. These can range anywhere from a sentence in length to a few paragraphs.
  • Scene or chapter scaffolding: Writers can provide a short summary, a list of plot points, or a few “beats” (key story moments), and the AI will generate a draft of a full scene or chapter based on that outline (Figure 6.
  • Story and worldbuilding tools: Writers can create detailed character profiles, track plot timelines, and maintain notes on fictional settings. Examples are Novelcrafter’s Codex: a manual, wiki-style database for structured character profiles, settings, and timelines, and NovelAI’s Lorebook: a repository for information about the story’s characters, settings, and other lore that is added to the AI’s context.
  • Model support: Some platforms allow users to choose between different AI models (like GPT-4, Claude, custom or fine-tuned models shared in the community) based on their writing preferences, as well as customize the model parameters. LAIKA featured ”brains”: AI writing partners trained on either your own writing, or public domain texts.
  • Prompt customization: Rather than starting from scratch each time, writers can save and reuse custom prompts and get support in creating prompts.
  • Editing and rewriting support: AI can help rewrite a passage to improve tone, make it more concise, or shift its style. Some tools even offer sentence-by-sentence suggestions or allow writers to highlight specific parts for revision.
  • Color-coded annotations: Some tools allow writers to mark which parts of the text were generated by AI, and which parts were written or edited by the author (Figure5).
Figure 5. NovelAI Interface, showing AI writing in white and human or human edited writing in blue. The panel on the right shows a few of the model settings.
[…]
A screenshot of NovelAI’s writing interface showing the tutorial chapter draft with prose about a king and his kingdom. A sidebar on the right shows model settings, including the AI model name: Kayra, AI Module (changes the style of the AI generated text): no module selected, Config Preset (changes the settings of the module: carefree, and Memory (the AI will better remember info placed here) in which there is the following text ”This is an epic fantasy story. My witch companion, Galena, and I are tasked to defeat a crystal dragon to protect the kingdom.”
Figure 6. A portion of Novelcrafter’s drafting interface. The SCENE BEAT feature allows writers to outline the scene for AI to generate. The toolbar shows options for the selected text, such as expanding, rephrasing, or shortening.
[…]
A screenshot of Novelcrafter’s drafting interface displays a draft of Chapter 1 with prose describing an underwater scene: ’The water vibrated beneath me, alive with a turquoise glow… A furry head popped up beside me, a wet nose nudging my arm.’ Below that is an empty text box labeled SCENE BEAT with a button labeled ”Generate Prose.” The text ”a wet nose nudging my arm” is selected and highlighted. Above that selection is a toolbar with text editing options and buttons labeled Expand, Rephrase, Shorten.
Appendix D Writers’ Values
We identify four core values that creative writers want to preserve while using AI: authenticity, ownership & control, creativity, and craftsmanship. The descriptions provided for each value aim to distill what writers mean when they refer to these concepts—while noting that participants often interpreted these values in different ways (Table 3).
Table 3. Writers’ values codes and example quotes.
@@ Appendix D Writers’ Values
Authenticity Importance of writing feeling that the writing comes from the author. (After AI generating a few lines) ”But beyond that, that’s when I start getting a little bit uncomfortable with how much am I losing myself in this writing process to something that’s not me authentically.” —P13 “Keeping it still sounding like me is, like, really important to me.” —P4
@@
Appendix E Participant Writing Session Setup
During the observed live-writing sessions, participants engaged in their writing practice as usual. Some worked on continuing in-progress pieces while others started on new projects. We summarize the writing task they worked on, how they used AI during that time, the specific AI used, and notes on how they set up their writing tools (Table 4).
Table 4. Live-writing session observed activity: summary of what each writer worked on during their writing sessions along with how AI was used.
ID Writing task Writing session activity AI used Notes on setup Writer’s View of AI
1 Chapter of thriller novel Using AI to do research and scene development. ChatGPT ChatGPT and writing document in separate tabs Tool
2 Chapter of romance novel Using AI to reword plot points, generate prose, and then manually editing and writing. Novelcrafter Workfing fully in Novelcrafter, keeping notes of custom prompts in Notion Tool, Director/Actor
3 Short story Using experimental custom GPT to take fragments of ideas to combine into new story. Editing existing section, using AI to flesh out motivations of characters. ChatGPT, custom GPTs ChatGPT and writing document in separate tabs Tool, Assistant
4 Personal essay Using AI to edit, reword, and get feedback. ChatGPT ChatGPT and writing document in separate tabs Tool, Editor
5 Sci-fi short story Using AI to generate inline continuations, choose from a branching interface, then manually continuing to write. code-davinci-002 Plugin that generates inline to the writing document (a branching text interface) Collaborator, Tool
6 Fanfiction Generate continuations inline and editing to correct logic, manually writing to direct dialogue and pacing. NovelAI Working fully in NovelAI Tool, Director/Actor, Ghostwriter
7 Chapter of fantasy novel Using AI to generate inline continuations. Editing and writing to guide AI toward specific story vision. NovelAI Working fully in NovelAI Director/Actor, Collaborator
8 Personal essay (starting new) Using AI to do research, and explore themes and styles. ChatGPT ChatGPT and writing document in separate tabs Tool, Collaborator
9 Poem (starting new) Using AI to generate a love poem, then manually editing. ChatGPT ChatGPT and writing document in separate tabs Tool
10 Romance novel Using AI to reword plot points, generate prose, and then manually editing and writing. Novelcrafter Writing in Novelcrafter, Perplexity for ideation Tool, Junior Writer, Assistant
11 Short story (starting new) Using AI to generate suggestions inline, chatting with personals for plot twists, new interpretations, character crafting. LAIKA Writing directly in Laika Tool, Muse
12 Chapter in new novel of a romance series Editing Novelcrafter generated prose in Claude to match writer vision. Claude, Novelcrafter Back and forth between Novelcrafter and Claude, keep track of ideas in Notion Co-writer, Assistant, Tool
13 Song lyrics (starting new) Using AI to generate themes and ideas, and restrcuture verses. ChatGPT ChatGPT and writing document in separate tabs Tool, Collaborator
14 Romance novel (starting new) Using AI to generate outline, edit outline, and generate prose. Rexy Using Rexy to write in Notion Tool, Director/Actor
15 Poem (starting new) Prompting ChatGPT to produce material on a theme, then collecting phrases to arrange into a poem. ChatGPT ChatGPT and writing document in separate tabs, collects by color Tool
16 Poem (starting new) Using AI to generate poem on a theme, then editing with further prompts (interpretation of the poem, new phrasings, new metaphors). ChatGPT ChatGPT and writing document in separate tabs Muse, Tool, Collaborator
17 Song lyrics (starting new) Using AI to generate lyrics from custom outline and prompt structure, then manually editing and rewriting. ChatGPT ChatGPT and writing document in separate tabs Tool
18 Poem Starting from found material (exisiting text) using AI to brainstorm ideas, then get feedback on drafts and advice on the genre. Claude, Copilot Claude/Copilot and writing document in separate tabs Tool, Editor
Appendix F Writers’ Relationships with AI
Table 5. How writers see AI’s role in the writing process.
@@ Appendix F Writers’ Relationships with AI
Appendix G Writers’ Specific Uses of AI
Table 6. Common AI use during observed writing sessions.